# A general framework for the generation of probabilistic socioeconomic scenarios: Quantification of national food security risk with application to the cases of Egypt and Ethiopia

### 07 Dec 2023 v1

@@ Abstract
In this work a general framework for providing detailed probabilistic socioeconomic scenarios as well as estimates concerning country-level food security risk is proposed. Our methodology builds on (a) the Bayesian probabilistic version of the world population model and (b) on the interdependencies of the minimum food requirements and the national food system capacities on key drivers, such as: population, income, natural resources, and other socioeconomic and climate indicators. Model uncertainty plays an important role in such endeavours. In this perspective, the concept of the recently developed convex risk measures which mitigate the model uncertainty effects, is employed for the development of a framework for assessment, in the context of food security. The proposed method provides predictions and evaluations for food security risk both within and across probabilistic scenarios at country level. Our methodology is illustrated through its implementation for the cases of Egypt and Ethiopia, for the time period 2019-2050, under the combined context of the Shared Socioeconomic Pathways (SSPs) and the Representative Concentration Pathways (RCPs).
@@
1 Introduction & Basic Concepts on Food Security
National food security is a key issue in sustainability studies, while a key factor in such studies is population Garnett (2013); Premanandh (2011). According to World Food Summit Summit (1996):
”Food security exists when all people, at all times, have physical and economic access to sufficient, safe and nutritious food that meets their dietary needs and food preferences for an active and healthy life”.
Under this definition, four different dimensions [^2] https://www.worldbank.org/en/topic/agriculture/brief/food-security-update/what-is-food-security are covered: (a) physical availability of food, (b) economic and physical access to food, (c) food utilization and (d) stability (sustainability) of the other three dimensions over time. In this work, we focus on the modelling, quantification and prediction of food security risk at country level, following a bottom-up probabilistic modelling approach relying mainly on a key factor that affects food security, i.e. population and its structure. Probabilistic population projections are generated under plausible socioeconomic scenarios and in particular under the so called Shared Socioeconomic Pathways (SSP) framework Lutz et al. (2017); Lutz et al. (2018), and then are incorporated to a global macroeconomic model for generating SSP-compatible projections for key drivers of national economies, like GDP, labour, etc. Then, these probabilistic modelling approaches, accompanied by standard econometric models for describing certain components that affect food security and taking into account the effect of climate drivers exogenously, are used to construct a decision support tool for assessing national food security risk. We insist on the term national food security since at a different scale, e.g. at household level, the situation could differ significantly [^3] Australia, Canada, UK and USA display high food security at national level but significant food insecurity at household level (see e.g. Loopstra (2018))). Therefore, in the perspective that food security risk is treated in this work, two main quantities of interest are considered as the key components: (a) the minimum food requirements of the population at a specific point in time (in terms of caloric content, rather than recommended dietary composition), and (b) the national food system capacity at the same time, i.e. the total quantity of the food (again in terms of caloric content) that is available to the country’s population. Note that finer aspects like dietary needs, food habits and poverty issues that may restrict access to food, are not directly treated in our approach. However, the proposed framework, can be extended to account for alternative modelling considerations, hence issues such as the above can be introduced if required. In what follows, by the term food requirements we will mean caloric requirements or intake.
Minimum food requirements for subsistence [^4] Note that the minimum caloric intake for bare survival does not necessarily coincide with the minimum calorie requirements that guarantee nutritional security depend in an inelastic way on population, as well as on its detailed age structure and each age group’s activity levels. On the other hand, food self-sufficiency at national level is mainly determined by the capacity of the country’s food system, indicated by the food amount (calories) available for human consumption as a result of the system functioning. This term is very accurately described in the definition provided by Food and Agriculture Organization (FAO) [^5] FAO (1999) https://www.fao.org/3/X3936E/X3936E00.htm:
“The concept of food self-sufficiency is generally taken to mean the extent to which a country can satisfy its food needs from its own domestic production”.
The capacity of the food system is importantly affected by the country’s economy, through the production system that involves planning of the economic sectors involved in food production and/or distribution, choice of crops, imports, exports, etc.
Reliable predictions for both minimum caloric requirements and food system capacity, in a long-time horizon, are very important in determining future food balance. Quantifying food security risk on a sufficiently long horizon, provides policy makers the luxury of adopting long-term measures, combining a portfolio of production restructuring policies, international trade treaties, adoption of scientific measures or modern technologies etc, which may efficiently alleviate the risk of future food security. Such trustworthy predictions will inevitably rely upon driver trends, and must be probabilistic in nature, i.e. providing the distribution of the relevant random factors, giving full information of the trends along with their validity, rather than point estimates which carry less information concerning the predictions. They should also take into account model uncertainty, which is inherent in these approaches, especially when long-term predictions are involved, because the effects of stochastic factors on these predictions may be at best partially known.
In the relevant literature, much attention has been given to the impacts of climate change in food security risk evaluations (see e.g. Schmidhuber and Tubiello (2007) and references therein). Discussions concerning the importance for the determination of measures and the actions to be taken in order to counter food security challenges due to climate change and extreme climate events are provided in Campbell et al. (2016) and Hasegawa et al. (2021). However, Hasegawa et al. (2018) argue that stringent climate mitigation policies would have a greater negative impact on global hunger and food consumption than the direct impacts of climate change. From the modelling perspective of food security, several directions are discussed in the literature introducing methodological frameworks (see e.g. Krishnamurthy et al. (2014)) and recognizing the need to better model key factors of food demand like demographics and economic growth due to the uncertainty that are subject to Valin et al. (2014). It seems that scenario-based approaches Hasegawa et al. (2015); Molotoks et al. (2021); Van Dijk et al. (2021) and attempts employing global models van Meijl et al. (2020) or integrating different models Müller et al. (2020), lead to more accurate assessments of food security risk, while simultaneously treating the model uncertainty issues. Although the majority of papers study the evaluation of the global food security risk, there are some works that study the matter on regional or country level Chen et al. (2021); Mainuddin and Kirby (2015) following the aforementioned approaches. In this work, a scenario-based approach is adopted for evaluating food security risk at country level, considering various socio-economic and climate scenarios under the probabilistic setting (i.e. each scenario is considered as a probability model which can produce a number of different trajectories) which to the best of our knowledge has not been yet proposed in the relevant literature. To this direction, the typical framework of the Shared Socioeconomic Pathways is revised to be compatible with this probabilistic setting, and then relevant modelling approaches are followed. Moreover, a food security risk index which is able to discount food security risk within and across the considered probabilistic socio-economic combined with climate scenarios is proposed.
@@ 1 Introduction & Basic Concepts on Food Security
More specifically, we provide a suitable framework for assessing food security risk at country level, following a bottom-up modelling approach, building on the probabilistic extension of the population model introduced in Raftery et al. (2012). The underlying population model relies on a Bayesian hierarchical modelling (BHM) procedure (see e.g. Gelman et al. (1995)) and provides detailed information on the future population of the world and its age structure by country under the probabilistic scenarios framework, i.e. for a certain time instant a distribution of the possible population counts is provided instead of a single point prediction. Our first contribution, is that by an appropriate modification of the structure of the population model, we are allowed to generate probabilistic population scenarios under the SSP framework, i.e. generating population trajectories of a priori known SSP status. Our second contribution is the incorporation of the generated population projections into a global macroeconomic model (in our case MaGE Fouré et al. (2013); Fontagné et al. (2022)) to generate SSP-compatible future scenarios for important drivers of economy like GDP, labour, etc. Finally, this probabilistic scenarios generation scheme is combined with climate scenarios, within the context of the Representative Concentrations Pathways (RCP) (see e.g. Meinshausen et al. (2020)), thus providing a complete framework for policies assessment for taking into account uncertainty on future predictions. Importantly, the probabilistic modelling framework adopted in this work, allows us to turn the qualitative narratives of the SSP/RCP scenarios into concrete quantitative scenarios, thus turning them into important quantitative policy and decision making tools (see Section 2.2).
The proposed framework is then implemented in quantifying food security risk at country level, for both between and across a selected set of combined socioeconomic and climate scenarios (SSP-RCP), under the aforementioned probabilistic setting. In this direction, the following modelling considerations are performed. First, employing the projected population structures (by age and gender), as well as the minimum required calories intake per age group and gender (determined by nutrition experts), estimations are provided for the minimum calories intake required for subsistence for the population at national level. Next, the national food system capacities under their current food system patterns are estimated, building models relying on basic socioeconomic drivers (population, income, labour) and climate factors (temperature, precipitation). These structural models of food system capacities are calibrated on past data and then, incorporating the population projections, as well as those of other related socioeconomic drivers, probabilistic projections on food system capacities are provided for a time period of interest (in our case 2019-2050). Combining the above modelling approaches, a new food security risk indicator is proposed. The advantages of the new index is that (a) it requires less detailed data for its calculation compared to well established food security risk indicators (please see the relevant discussion at the beginning of Section 3), (b) it is more affordable in terms of computational cost, (c) it is able to project the food security situation in the future, and (c) it takes into account natural resources, such as for example the pressure on national water sources. The construction of this index relies on the concept of the convex risk measures (see e.g. Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)) allowing for mitigating successfully the effects of model uncertainty and providing robust risk estimates either within a certain probabilistic scenario or across different scenarios. The latter risk estimates rely on the concept of Fréchet utilities or risk measures (Papayiannis and Yannacopoulos (2018), Petracou et al. (2022)) and are crucial, especially for long-term effects and policies, that have to be predicted and implemented long before the actual scenario that materializes has been fully clarified.
The paper is organized in the following manner: In Section 2 the probabilistic extension of the population model and its modification so that it becomes compatible with the SSP framework and the set of socio-economic and climate scenarios that are considered throughout the paper are introduced. Next, in Section 3 the modelling approaches and considerations for the minimum caloric intake and the food system capacity at country level are presented. Moreover, the new food security risk index is introduced, accompanied by a discussion and comparisons to relevant well established indicators in the literature. Finally, in Section LABEL:sec-3, the proposed approach is implemented for the cases of Egypt and Ethiopia using as training data records from the period 1990-2018 and providing projections till the year 2050 both for within and across the SSP-RCP scenarios framework.
2 A general probabilistic socioeconomic modelling framework for food security
@@ 2.1 The main probabilistic population model
Population growth and evolution of the population’s age structure are key factors driving many socioeconomic indices, including economic growth, production, environmental issues, food and water demand etc. In this respect, demographics must be the starting point for any socioeconomic modelling study. In this section, some key results concerning scenario development for future population growth are presented. The state of the art model concerning world population is the probabilistic model proposed by Raftery and coworkers Raftery et al. (2012). This model takes into account the inherent uncertainty of the phenomenon and its effects on future population projections using a Bayesian hierarchical model (see e.g. Congdon (2010); Gelman et al. (1995)). The model is based on the natural evolution of the population phenomenon as characterized by the standard (deterministic) model employed by the United Nations (UN),
(1) $P_{c}(t)-P_{c}(t-1)=B_{c}(t)-D_{c}(t)+M_{c}(t)$
where $P_{c}(t)$ denotes the population of country $c$ at time $t$ (corresponding either to a single year or a 5-year period), $B_{c}(t)$ stands for the number of births (which depends on the total fertility rate of the country), $D_{c}(t)$ denotes the number of deaths (which depends on the life expectancy) and $M_{c}(t)$ measures the net international migration. Uncertainty is introduced to the population model since its main components (fecundity, mortality, migration) are subject to random factors that cannot be sufficiently modelled. The Bayesian approach proposed in Raftery et al. (2012) captures uncertainty on each one of the major components of population through the construction/introduction of distinct hierarchical models for important components such as fertility, mortality and migration [^6] Note that only international migration is considered and not internal migration phenomena like urbanization, and then propagating uncertainty to the output of model (1), i.e. providing probabilistic estimates either for $P_{c}(t)$ or its breakdown into age groups and sex at various future times $t$. Clearly, uncertainty becomes higher as the time progresses. Based on an extensive database of past world population data (recorded population pyramids, fertility rates, mortality rates, etc), the fundamental law (1), and the principles of Bayesian statistics, the probabilistic features of the uncertainty factors driving the population fluctuations are recovered. Then, using this information, the fundamental law (1) is iterated forward and used to obtain estimates for the future evolution of the quantities of interest. The estimates incorporate in a dynamically consistent fashion the effects of ambiguity as documented at least from the past data, and thus provide uncertainty consistent predictions for the future.
@@ 2.2 Shared Socioeconomic Pathways, Population and GDP Projections
The concept of scenario making concerning future events has infiltrated environmental economics and has become a fundamental tool in the analysis of potential future outcomes. An important class of scenarios used frequently in analyses are the Shared Socioeconomic Pathways O’Neill et al. (2014), which set certain plausible assumptions concerning the evolution for key socioeconomic drivers (such as population growth or fertility) for certain parts of the world. The possible states of the world are divided into five qualitative scenarios (rapid, medium, stalled, inequality and development) according to the levels of specific demographic characteristics and specifically, fertility, life expectancy (or mortality), migration and education. Since the adopted population projection method does not take into account the education levels, this factor is omitted for the purposes of this work.
These narratives are phrased in a qualitative fashion (please see Table 1), so they have to be transcribed to quantitative terms if they are to be used constructively in quantitative models. For instance, under the SSP1 scenario presented in Table 1 the evolution of life expectancy in high fertility countries is characterised as high, with the obvious question arising being: “How high is high? Which figure for the life expectancy in a particular country can safely fit (i.e. with a particular probability) to the qualitative characterization high? An important contribution of our work, and a key point in our approach is the transcription of the SSPs in a quantitative, probabilistic setting. This allows us to provide the population projections (appropriately quantified in terms of probability distributions or sample paths for the quantities of interest) under each one of these pathways.
@@
SSP scenarios may differ for each country, depending on its grouping as (a) high fertility country (HiFert), (b) low fertility country (LoFert) or (c) Rich-OECD country [^8] http://www.oecd.org/about/members-and-partners/. Specifications of the various SSPs with respect to the country groupings, as described in Lutz et al. (2017); Lutz et al. (2018), are illustrated in Table 1. As already mentioned, serious drawback of defining the scenarios in this fashion, is the difficulty to adopt a common line on what is meant by low, medium and high for the various quantities of interest. This is of paramount importance in quantitative modelling, upon which policy making will be based. For instance, assume that we need to quantify the low fertility sub-scenario for both a LoFert and a HiFert country. According to the UN fertility modelling the value 2.1 is considered as the threshold value under which the population of a country declines and above which the population increases.
If we are planning to generate population scenarios for a relative short time period (e.g. 20-30 years from now), according to the aforementioned threshold for TFR, we might not be able to generate all possible TFR sub-scenarios for all countries, due to the different dynamics of the collected data. In this perspective, for a LoFert country an SSP scenario that includes ”high TFR” may not be feasible, while not feasible might be also for a HiFert country an SPP scenario that includes ”low TFR”. Therefore, this universal determination for high-low cases leads to the problematic situation that a policy maker might not be able to create all the SSP scenarios for all countries at any time horizon, a fact that can cause problems in environmental modelling. However, as shown in this section, the probabilistic approach to modeling (such as for example the population model described in Section 2.1) can be nicely integrated with SSP qualitative scenarios turning them into well defined quantitative tools. This, provides a concrete and realistic framework for scenario building, in which the various sub-scenarios (Low, Medium and High) are endogenously and consistently determined by the evolutionary dynamics of the system under study.
The quantification of the SSP scenarios can be effected as follows: Using the probabilistic population model presented in Section 2.1, samples of trajectories for the possible realizations of the evolution over time for the fertility and life expectancy components for each country of interest are generated. These samples are used to obtain the evolution of the probability distribution of these quantities over time, and the relevant quantiles of these distributions at specific time instances are then used to determine the quantitative thresholds that characterise the sub-scenarios Low, Medium and High, as mentioned in Table 1. For instance, if we are interested in determining the range of the three different sub-scenarios (Low, Medium and High) for the TFR of a certain country $c$ up to a certain time horizon $T$ (e.g. $T=2050$), we operate as follows: Employing the distribution of the generated trajectories at the time instant $T$ for the particular country of interest $c$, we separate the induced distribution in three equal parts using the 33% and 66% quantiles, and then use these values as lower and upper thresholds to distinguish the sample of trajectories to the three different sub-scenarios. According to this rule, a trajectory is assigned to the high sub-scenario if at the terminal year 2050 the observation for the corresponding quantity lies on the top 66% of the empirical distribution (i.e. if for trajectory $j$ holds that $TFR_{c,2050,j}\geq q_{TFR,c,2050}(0.66)$ with $q_{TFR,c,2050}(\cdot)$ denoting the induced from the generated sample quantile function for TFR at year $2050$ for the country $c$). The classification of trajectories to the other sub-scenarios is performed in a similar manner.

### 14 Dec 2023 v2

# Probabilistic Scenario-Based Assessment of National Food Security Risks with Application to Egypt and Ethiopia

@@ Abstract
This study presents a novel approach to assessing food security risks at the national level, employing a probabilistic scenario-based framework that integrates both Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP). This innovative method allows each scenario, encompassing socio-economic and climate factors, to be treated as a model capable of generating diverse trajectories. This approach offers a more dynamic understanding of food security risks under varying future conditions. The paper details the methodologies employed, showcasing their applicability through a focused analysis of food security challenges in Egypt and Ethiopia, and underscores the importance of considering a spectrum of socio-economic and climatic factors in national food security assessments.
@@
1 Introduction
In the era of rapid global change, understanding the intricate relationship between population dynamics, climate change, and food security has become more critical than ever. Evolving demographic trends, coupled with climatic shifts, are shaping the future landscape of food availability and access, calling for a comprehensive analysis that delves into the multifaceted nature of food security, acknowledging that it is not solely an issue of production but also one deeply intertwined with socioeconomic and environmental factors. The global population is projected to reach nearly 10 billion by 2050, introducing profound challenges and demands on the food supply systems. Population growth, urbanization, and demographic transitions significantly influence food demand patterns, dietary preferences, and the overall caloric requirements of nations. This demographic shift, while a central element, does not exist in isolation. It interacts with and is influenced by broader socioeconomic changes, including economic development, technological advancements, and policy frameworks. These factors collectively dictate the availability, accessibility, and utilization of food resources, thereby shaping national food security profiles. In addition, climate change emerges as a formidable force, altering the fundamentals of food production. Shifts in temperature, precipitation patterns, and increased frequency of extreme weather events have profound implications for agricultural productivity, food prices, and ultimately, the stability of food supply chains. The growing body of literature underscores the profound impact of climatic shifts on food systems worldwide. Discussions focus on the need for effective measures to tackle food security challenges posed by climate change and extreme events. However, there is a view that stringent climate mitigation policies might have more adverse effects than climate change itself on global hunger and food consumption. It is crucial to understand how climatic factors interact with population dynamics to anticipate future food security challenges. Methodological frameworks, scenario-based approaches, and global model applications are recognized for their effectiveness in providing accurate food security risk assessments, considering uncertainties in factors like demographics and economic growth. While much research assesses global food security risks, national food security and the role of population dynamics are also key concerns, with some studies focusing on regional or national levels.
In this study, our focus is on modeling, quantifying, and predicting food security risks at the national level. We adopt a bottom-up probabilistic modeling approach, emphasizing a critical determinant of food security: population and its structure. We generate probabilistic population projections within the framework of plausible socioeconomic scenarios, particularly under the Shared Socioeconomic Pathways (SSP) framework Lutz et al. (2017); Lutz et al. (2018). These projections are then integrated into a global macroeconomic model to produce SSP-compatible forecasts for essential drivers of national economies, such as GDP and labor. Utilizing these probabilistic models, we aim to contribute towards a more holistic assessment of national food security risks. Our approach offers a multi-dimensional view, considering how socio-economic development paths might mitigate or exacerbate the impacts of climate change on food security. Given the complexities outlined, we recognize the need to integrate both socioeconomic and climatic considerations adopting a comprehensive framework that blends the socioeconomic scenarios with climate scenarios under the Representative Concentration Pathways (RCP) framework Meinshausen et al. (2020). This integration is crucial as it allows us to capture the interplay between human socio-economic development and climatic changes, offering a more holistic understanding of food security risks.
Accurate long-term forecasts of caloric requirements and total caloric content of food available to the population are crucial in assessing future food balance. By quantifying the disparities between these two metrics over an extended time horizon, policymakers are afforded the opportunity to implement long-term strategies to avoid food risks. These strategies can encompass a diverse portfolio, including restructuring agricultural production, negotiating international trade agreements, and adopting scientific advances or modern technologies. Such measures can effectively mitigate potential future food security risks. For these predictions to be reliable, they must be grounded in probabilistic modeling, which provides the full distribution of relevant variables rather than mere point estimates. This approach offers a more complete picture of future trends and their probabilities, adding a depth of understanding that point estimates lack. Additionally, it is vital to incorporate model uncertainty into these predictions. This aspect is particularly critical in long-term forecasting, where the impact of stochastic factors can only be partially anticipated. Acknowledging and accounting for this uncertainty ensures a more robust and realistic assessment of potential future scenarios in food security.
@@ 1 Introduction
To this end, we introduce a food security risk index designed to effectively quantify disparities between caloric requirements and total available caloric content of food within and across the devised socio-economic and climate scenarios. Our study offers a specialized framework for assessing food security risks at the national level, employing a bottom-up modeling approach. This builds upon the probabilistic extension of the population model as outlined in Raftery et al. (2012). The model, based on a Bayesian hierarchical modeling procedure (BHM) as seen in Gelman et al. (1995), provides detailed projections of the world’s future population and its age structure by country, under various probabilistic scenarios. Unlike traditional models that offer a single point prediction, our approach yields a distribution of possible population counts for a specific time instant. Our first major contribution is the modification of this population model structure, allowing us to generate probabilistic population scenarios within the SSP framework. This enables the creation of population trajectories with predefined SSP statuses. The second key contribution is integrating these population projections with a global macroeconomic model, specifically MaGE Fouré et al. (2013); Fontagné et al. (2022), to produce SSP-compatible projections for vital economic drivers such as GDP and labor. Furthermore, we blend this probabilistic scenario generation scheme with climate scenarios under the Representative Concentration Pathways (RCP) framework Meinshausen et al. (2020). This comprehensive approach provides a robust framework for policy assessment, considering uncertainties in future predictions. Crucially, our probabilistic modeling framework transforms the qualitative narratives of SSP/RCP scenarios into actionable quantitative scenarios, making them valuable tools for quantitative policy and decision-making (refer to Section 2.2 for details).
Our modeling process encompasses several steps. Initially, using the projected population structures (by age and gender), and minimum caloric intake requirements per age group and gender (as determined by nutrition experts), we estimate the minimum caloric intake necessary for subsistence at the national level. Subsequently, we assess the capacities of national food systems under their existing patterns, forming models based on fundamental socioeconomic drivers (population, income, labor) and climate factors (temperature, precipitation). These models, calibrated with historical data, are then combined with projected population and other socioeconomic drivers to provide probabilistic forecasts of food system capacities up to 2050. Building on these quantities, we introduce a novel food security risk indicator. This index offers several advantages: (a) it requires less detailed data for computation compared to established food security risk indicators (discussed in Section 3), (b) it is computationally efficient, (c) it projects future food security scenarios, and (d) it incorporates the impact on natural resources, such as national water sources. The foundation of this index is the concept of convex risk measures (referenced in Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)), which effectively mitigate the effects of model uncertainty, providing robust risk assessments within and across various probabilistic scenarios. These assessments are based on the concept of Fréchet utilities or risk measures (Papayiannis and Yannacopoulos (2018), Petracou et al. (2022)), which are particularly valuable for long-term policy planning and implementation, crucial in scenarios where the exact future circumstances have yet to be fully determined.
The paper is structured as follows: Section 2 details the probabilistic extension of the population model and its adaptation to align with the Shared Socioeconomic Pathways (SSP) framework. This section also introduces the set of socio-economic and climate scenarios that are central to our analysis. Subsequently, Section 3 delves into the modeling approaches and considerations for determining the minimum caloric intake and evaluating the capacity of food systems at the national level. In addition, this section presents our newly developed food security risk index, along with a comprehensive discussion and comparative analysis with established indicators in the field. Finally, Section 4 applies our proposed approach to specific case studies: Egypt and Ethiopia. Utilizing data from 1990 to 2018, we offer projections up to the year 2050, examining both within and across various SSP-RCP scenario frameworks. This practical implementation demonstrates the applicability and effectiveness of our methodology in real-world settings.
2 Socio-Economic and Climatic Scenario Modeling
From a modeling perspective, the literature explores various methodological approaches to food security, as outlined in Krishnamurthy et al. (2014), and emphasizes the importance of accurately modeling crucial food demand factors such as demographics and economic growth, given their inherent uncertainties Valin et al. (2014). Scenario-based approaches Hasegawa et al. (2015); Molotoks et al. (2021); Van Dijk et al. (2021) and global model applications van Meijl et al. (2020), as well as integrative modeling efforts Müller et al. (2020), are recognized for their effectiveness in providing more accurate food security risk assessments while addressing model uncertainty. While the majority of such research focuses on assessing global food security risks, national food security is a pivotal concern in sustainability studies, with population dynamics playing a crucial role in this context Garnett (2013); Premanandh (2011). To this respect, other studies that explore food risks at regional or national levels are Chen et al. (2021); Mainuddin and Kirby (2015). In this study, we implement a scenario-based approach for assessing food security risk at a national level, utilizing a diverse array of socio-economic and climate scenarios within a probabilistic framework. This method, where each scenario is treated as a model capable of generating multiple trajectories, is a novel approach in the field. We have adapted the standard framework of the Shared Socioeconomic Pathways to align with this probabilistic setting.
@@ 2.1 The main probabilistic population model
The dynamics of population growth and the evolution of age structures are fundamental in driving numerous socioeconomic indices, including economic growth, production capacities, environmental challenges, and the demand for food and water. Consequently, demographic analysis is an essential starting point for any socioeconomic modeling study. In this section, we present some key findings related to the development of scenarios for future population growth. The leading model in this domain is the probabilistic approach proposed by Raftery and colleagues Raftery et al. (2012), which addresses the inherent uncertainties in demographic forecasting. This model employs a Bayesian hierarchical approach (referenced in Congdon (2010); Gelman et al. (1995)) to account for the uncertainties and variability affecting future population projections. The model is based on the natural evolution of the population phenomenon as characterized by the standard (deterministic) model employed by the United Nations (UN),
(1) $P_{c,t}-P_{c,t-1}=B_{c,t}-D_{c,t}+M_{c,t}$
where $P_{c,t}$ denotes the population of country $c$ at time $t$ (corresponding either to a single year or a 5-year period), $B_{c,t}$ stands for the number of births (which depends on the total fertility rate of the country), $D_{c,t}$ denotes the number of deaths (which depends on the life expectancy) and $M_{c,t}$ measures the net international migration. Uncertainty is introduced to the population model since its main components (fecundity, mortality, migration) are subject to random factors that cannot be sufficiently modelled. The Bayesian approach proposed in Raftery et al. (2012) captures uncertainty on each one of the major components of population through the construction/introduction of distinct hierarchical models for important components such as fertility, mortality and migration [^2] Note that only international migration is considered and not internal migration phenomena like urbanization, and then propagating uncertainty to the output of model (1), i.e. providing probabilistic estimates either for $P_{c,t}$ or its breakdown into age groups and sex at various future times $t$. Clearly, uncertainty becomes higher as the time progresses. Based on an extensive database of past world population data (recorded population pyramids, fertility rates, mortality rates, etc), the fundamental law (1), and the principles of Bayesian statistics, the probabilistic features of the uncertainty factors driving the population fluctuations are recovered. Then, using this information, the fundamental law (1) is iterated forward and used to obtain estimates for the future evolution of the quantities of interest. The estimates incorporate in a dynamically consistent fashion the effects of ambiguity as documented at least from the past data, and thus provide uncertainty consistent predictions for the future.
@@ 2.2 Shared Socioeconomic Pathways, Population and GDP Projections
The practice of developing scenarios for future events has become integral in environmental economics, serving as a key tool for analyzing potential outcomes. A significant category of these scenarios is the Shared Socioeconomic Pathways (SSPs) O’Neill et al. (2014), which outline plausible developments for critical socioeconomic drivers such as population growth and fertility in various regions of the world. These SSPs categorize potential global futures into five qualitative narratives—rapid, medium, stalled, inequality, and development—based on specific demographic characteristics like fertility, life expectancy (or mortality), migration, and, although not considered in our method, education.
These SSP narratives, presented qualitatively in Table 1, require translation into quantitative terms for effective incorporation into quantitative models. For example, the SSP1 scenario describes the evolution of life expectancy in high-fertility countries as ‘high’. This leads to the question: What quantitatively constitutes ‘high’ life expectancy in a specific country? How can we assign a definitive figure to this qualitative descriptor with an associated probability?’ A pivotal aspect of our work is this translation of SSP narratives into a quantitative, probabilistic framework. This adaptation enables us to provide population projections that are quantified in terms of probability distributions or sample paths for the relevant demographics under each SSP pathway.
@@
The SSP scenarios are tailored to each country, classified as either a high fertility country (HiFert), a low fertility country (LoFert), or a Rich-OECD country, with specific SSP characteristics based on these groupings Lutz et al. (2017); Lutz et al. (2018), as detailed in Table 1. A significant issue with defining scenarios in this manner is the challenge in universally applying qualitative descriptors like low, medium, and high to different quantities of interest. This precision is crucial in quantitative modeling, which forms the foundation of policy-making.
Consider, for instance, the task of quantifying the ’low fertility’ sub-scenario for both LoFert and HiFert countries. The UN’s fertility modeling uses a threshold value of 2.1, interpreting values below this as indicative of a declining population and values above it as a growing population. However, when generating population scenarios for a relatively short term (say, 20-30 years), this fixed threshold may not accommodate all possible Total Fertility Rate (TFR) sub-scenarios due to varying data dynamics. For a LoFert country, a ’high TFR’ case might be unrealistic, just as a ’low TFR’ case might be improbable for a HiFert country. This rigid application of high-low thresholds can lead to a problematic situation where policymakers are unable to create comprehensive SSP scenarios for all countries across all time horizons, posing significant challenges in environmental modeling.
However, as we demonstrate in this section, adopting a probabilistic approach to modeling (like the population model described in Section 2.1) can effectively integrate with the qualitative SSP scenarios to transform them into well-defined quantitative tools. This approach offers a realistic and concrete framework for scenario building, where the various sub-scenarios (Low, Medium, and High) are endogenously and consistently determined by the system’s evolving dynamics, rather than relying on rigid, deterministic cut-offs. This methodology enables more flexible and accurate scenario construction, crucial for effective environmental modeling and policy planning. The quantification of SSP scenarios, as an alternative to deterministic cut-offs, can be effectively achieved through the probabilistic population model discussed in Section 2.1. This model allows for the generation of sample trajectories representing possible evolutions over time for fertility and life expectancy in each country of interest. By analyzing these samples, we can trace the evolution of the probability distribution of these demographic components over time. Subsequently, specific quantiles of these distributions at designated time points are utilized to establish the quantitative thresholds that define the Low, Medium, and High sub-scenarios, as referenced in Table 1. For example, to determine the range for the three TFR sub-scenarios (Low, Medium, High) for a country $c$ by a certain year $T$ (e.g., $T=2050$), we use the distribution of the projected trajectories at $T$. These projections are divided into three equal segments based on the 33% and 66% quantiles. These quantile values then act as the boundaries differentiating the trajectory samples into the respective sub-scenarios. A trajectory is classified as high if, in the year 2050, its corresponding measure falls within the top 33% of the empirical distribution. Specifically, for a trajectory $j$, it would be categorized as ’high’ if $TFR_{c,2050,j}\geq q_{TFR,c,2050}(0.66)$, where $q_{TFR,c,2050}(\cdot)$ is the quantile function derived from the sample for TFR in 2050 for country $c$. The allocation of trajectories to the other two sub-scenarios, Low and Medium, follows a similar methodology.